Please Help!! Find an equation of the line that satisfies the given conditions
Through (-2,4) slope is -1

Answer by CharlesG2(834)

Please Help!! Find an equation of the line that satisfies the given conditions
Through (-2,4) slope is -1

equation of a line in slope-intercept form is y = mx + b, where m is the slope, and b is the y-intercept (vertical-intercept or point (0,b))

m = -1, point is (-2,4), solve for b, then plug in b and get equation of line

y = -1 * x + b = -x + b
4 = -(-2) + b
4 = 2 + b
2 = b

y = -x + 2